### v1

# GridRoute: A Benchmark for LLM-Based Route Planning with Cardinal Movement in Grid Environments

Abstract
Recent advancements in Large Language Models (LLMs) have demonstrated their potential in planning and reasoning tasks, offering a flexible alternative to classical pathfinding algorithms. However, most existing studies focus on LLMs’ independent reasoning capabilities and overlook the potential synergy between LLMs and traditional algorithms. To fill this gap, we propose a comprehensive evaluation benchmark GridRoute to assess how LLMs can take advantage of traditional algorithms. We also propose a novel hybrid prompting technique called Algorithm of Thought (AoT), which introduces traditional algorithms’ guidance into prompting. Our benchmark evaluates six LLMs ranging from 7B to 72B parameters across various map sizes, assessing their performance in correctness, optimality, and efficiency in grid environments with varying sizes. Our results show that AoT significantly boosts performance across all model sizes, particularly in larger or more complex environments, suggesting a promising approach to addressing path planning challenges.111Our code is open-sourced at https://github.com/LinChance/GridRoute.
GridRoute: A Benchmark for LLM-Based Route Planning with Cardinal Movement in Grid Environments
Kechen Li1, Yaotian Tao2, Ximing Wen3, Quanwei Sun2, Zifei Gong4, Chang Xu2, Xizhe Zhang2, Tianbo Ji1,††thanks: Corresponding author 1School of Transportation and Civil Engineering, Nantong University, China 2HongKong Linchance Intelligent Technology Co., Ltd. (LinChance), China 3College of Computing and Informatics, Drexel University, USA 4University of New South Wales, Australia kechen.li@stmail.ntu.edu.cn, {taoyaotian,quanweisun,xuchang,zhangxizhe}@linchance.com xw384@drexel.edu, z3224950@zmail.unsw.edu.au, jitianbo@ntu.edu.cn
1 Introduction
Pathfinding in grid-based environments has long been dominated by classical algorithms such as A* Hart et al. (1968), Dijkstra (Dijkstra, 1959), and DFS, which rely on explicit search heuristics and guarantee optimality under predefined conditions. However, the recent rise of Large Language Models (LLMs) has introduced an alternative paradigm, everaging their implicit reasoning capabilities and contextual adaptation.
Specialized benchmarks, such as Path Planning from Natural Language (PPNL) Aghzal et al. (2023) explored LLMs’ ability to perform end-to-end navigation while adhering to movement constraints and avoiding obstacles. However, they only explored the independent exploring ability of LLMs using prompting strategies such as Chain-of-Thought (CoT) Wei et al. (2023), ReAct Yao et al. (2023). LLM-A* Meng (2025), on the other hand, tried to integrate LLM’s ability with guidance from traditional pathfinding algorithms, but they only use LLM in one of the steps, which is to generate intermediate target lists, and still rely on A* to generate the final solution. In this way, there still lacks research exploring how traditional algorithms could help LLM-generated route planning.
To bridge this gap, we propose a comprehensive benchmark GridRoute to systematically evaluate the role of traditional algorithmic guidance in enhancing the pathfinding capabilities of LLMs. Our benchmark is designed around three core prompting paradigms: (1) independent prompting, (2) Algorithm of Thought(AoT) prompt, where traditional algorithms such as DFS, A*, and Dijkstra provide reasoning guidance; and (3) hybrid prompting, which leverages algorithm-generated trajectories or decisions within both direct-answer and reasoning-based prompts. This design facilitates a comprehensive comparison of traditional algorithmic guidance versus independent prompting approaches, and further allows for in-depth analysis of variation within algorithm-guided prompts.
To assess scalability and generalization, we construct grid-based environments of three different sizes with increasing complexity. We evaluate the performance of six LLMs, ranging from 7B to 72B parameters, including models from multiple families, such as GPT-4 OpenAI et al. (2024), DeepSeek DeepSeek-AI et al. (2024), and QWEN Qwen et al. (2025). Model performance is assessed using metrics that capture correctness, optimality, and efficiency—-such as format compliance, path validity, solution quality, error from the optimal path, and runtime. This multifaceted evaluation offers a rigorous analysis of how algorithmic guidance affects LLM pathfinding, both in terms of accuracy and efficiency.
Through this benchmark, we aim to answer several key questions: To what extent can LLMs benefit from classical search heuristics? How do the different shot strategies affect the traditional algorithm’s guidance for LLM? Do algorithm-guided prompt performance remain indispensable regardless of model scale and map size? Our findings suggest that algorithm-guided prompts achieve better results than vanilla prompts and comparable performance compared with the CoT prompt. While larger models tend to perform better under independent prompting, algorithm-guided strategies significantly boost performance across all model sizes, particularly in larger or more complex environments. Moreover, integrating few-shot learning could further improve the performance of algorithm algorithm-guided prompt. These results highlight the complementary strengths of LLMs and classical algorithms and point toward promising directions for hybrid neuro-symbolic planning systems.
In conclusion, this study makes three major contributions:
We propose GridRoute, a comprehensive benchmark designed to evaluate LLMs’ route planning abilities in grid-based environments, comparing independent prompting and algorithm-guided strategies.
We introduce AoT prompting, a novel framework that guides LLMs using classical pathfinding algorithms (e.g., A*, Dijkstra, DFS) by embedding their reasoning principles directly into the prompt.
We perform large-scale experiments across multiple LLMs and environment sizes, showing that AoT consistently improves planning performance, especially in complex settings, and offers complementary benefits to traditional CoT prompting.
2 Related Work
2.1 Grid-based route planning
Grid-based route planning is central to robotics and embedded systems, especially in warehouse automation, where robots navigate structured indoor environments such as aisles and shelves. These environments are typically modeled as 2D grids with discrete free or occupied cells, offering memory and computational efficiency suitable for embedded platforms. Algorithms such as A* (Hart et al., 1968), Dijkstra’s Dijkstra (1959), and variants like Jump Point Search (Harabor and Grastien, 2014) are widely used for computing collision-free, cost-effective paths. Many systems adopt 4-directional movement (up, down, left, right) to reflect real-world layouts and simplify pathfinding. This constraint, along with lightweight planning algorithms, enables real-time decision-making under tight power and timing constraints, making grid-based navigation viable on low-power processors.
2.2 Route planning with LLM
Recent advances in large language models (LLMs) have showcased their emerging capabilities in reasoning, planning, and problem-solving. Unlike classical algorithms, LLMs leverage implicit knowledge and contextual understanding, allowing them to generate flexible solutions in unfamiliar scenarios (Wang et al., 2023; Li et al., 2025; Wen, 2025). Early studies have demonstrated that LLMs can produce plausible paths or high-level plans in grid-based environments when provided with textual descriptions of the environment, along with specified start and goal positions (Valmeekam et al., 2023; Latif, 2024). However, these outputs are often inconsistent or suboptimal, particularly as the complexity of the environment increases.
To address these limitations, recent research has introduced structured reasoning techniques to improve LLM-based planning. For example, the PPNL benchmark investigates the use of Chain-of-Thought (CoT) and ReAct prompting to encourage step-by-step reasoning in route planning tasks. While these methods improve performance over naive prompting, they still do not fully leverage the systematic search capabilities of classical algorithms. In parallel, an emerging line of work explores hybrid approaches that integrate LLMs with traditional pathfinding algorithms such as A* and Dijkstra (Liu et al., 2023). In these settings, LLMs may generate sub-goals for symbolic planners to solve or invoke classical algorithms as external tools within a CoT-style reasoning framework.
Building on this line of work, our study systematically investigates how classical algorithms can be used to guide LLM-based route planners. We focus on grid-based environments, where structured search procedures are well-defined, computationally efficient, and particularly suited for real-time or embedded applications.
3 GridRoute Benchmark
This section introduces the design of the GridRoute benchmark, including data generation, evaluation metrics, and error analysis. We first describe how the task environments are constructed and presented to LLMs, followed by the metrics used to assess performance. Finally, we analyze common failure modes to better understand the limitations of LLMs in route planning tasks.
3.1 Data Generation
GridRoute is proposed to evaluate the performance of LLMs on the task of automated route planning, and its overall structure is illustrated in Figure 1. In GridRoute, LLMs are required to discover a valid path from the origin to the destination in a randomly generated grid environment, while avoiding a set of non-overlapping square obstacles of size . The movement of the agent is constrained to four cardinal directions: up, down, left, and right.
Environment Setting To give a comprehensive evaluation of different prompting strategies, we employ three sets of map configurations in our experiments with spatial layouts of varying complexity: , , and . For each generated environment, five pairs of origin and destination points are randomly sampled. To ensure task difficulty and meaningful planning, each pair must satisfy a minimum Euclidean distance constraint of at least 30% of the grid’s diagonal length. In addition, to ensure path feasibility, we verify connectivity using a four-directional Breadth-First Search (BFS) algorithm.
Task Description All environments and tasks are presented to LLMs in natural language. To this end, we design standardized templates to verbalize structured task information, including grid size, obstacle locations, and origin/destination coordinates, into consistent and interpretable natural language prompts. For evaluation, we employ a constrained diagonal Dijkstra algorithm to compute the shortest paths for all task instances, yielding both the reference trajectory and the corresponding optimal path length.
Although this study focuses on the three predefined configurations, the GridRoute framework is inherently extensible and supports flexible adjustments to map size, obstacle density, and obstacle shapes. This design encourages future exploration of LLMs’ generalization capabilities in larger and more complex planning scenarios.
3.2 Evaluation Metrics
We evaluate the benchmark from six perspectives to assess the accuracy, quality, and efficiency of LLM-generated routes, as detailed below:
Compliance Ratio (CR): It indicates whether an LLM can produce a specific formatted output as instructed, namely a route in this case. A continuous route which starts from the origin and ends at the destination can be deemed “compliant”. It can be simply computed as where and are the numbers of compliant routes and test cases, respectively.
Feasibility Ratio (FR): It shows whether an LLM can generate a feasible route, where a “feasible” route is “compliant” and avoid any obstacle in it. FR is computed as .
Optimal Ratio (OR): It is used to assess whether a feasible LLM-generated route is optimal compared to the Dijkstra ground truth according to the length. OR is computed as .
Geometric Mean (GM): It measures the similarity between LLM-generated routes and reference optimal routes. It is computed as where returns the length of a given route, and are the LLM-generated route and the Dijkstra ground-truth route, respectively. Higher LLM performance is associated with GM approaching 1.
Mean Square Error (MSE): It measures the difference between candidate routes and optimal routes, computed as . MSE closer to 0 indicates a higher performance.
Run Time (RT(s)):It records the average time for LLMs to generate paths of different sizes.
3.3 Route Failure Types
As part of the GridRoute benchmark, we further categorize the generated failure cases into five representative error types, including:
Invalid Step Distance: Indicates the presence of non-unit-length steps (e.g., diagonal moves) in the path, violating grid-based movement constraints.
Path Through Obstacle: Occurs when the generated path intersects predefined obstacle regions, failing to satisfy collision-free requirements.
Out of Bounds: Reflects paths exceeding map boundaries, revealing the LLMs’ lack of spatial awareness regarding environmental constraints.
Empty Path: Denotes failure cases where the LLM produces no path output.
Start/End Mismatch: Identifies paths that either deviate from the specified origin or fail to reach the designated destination.
4 Experiments
4.1 Experiments Setup
In our experiments, we conducted a comprehensive evaluation involving the Qwen2.5 series, LLaMA3.1-70B (Grattafiori et al., 2024), DeepSeek-V3, and GPT-4 Turbo. The Qwen2.5 series includes models of various parameter scales, specifically 7B, 32B, and 72B while DeepSeek-V3 has 671 billion parameters.
4.2 Prompt Design
To evaluate how the guidance of traditional algorithms influences the ability to plan the route of LLMs, we experiment with mainly three categories of prompting strategies, with examples of each prompt shown in the appendix A.
4.2.1 Independent Route Planning Prompt
In this part, we use three types of prompts to evaluate the independent planning capabilities of LLMs: vanilla instruction-based prompting, CoT prompting, and Few-shot Learning prompting.
Vanilla: it is the basic prompt which provides straightforward instructions for LLMs to complete the path planning task, by which LLMs are simply asked to generate a continuous, obstacle-free route given the map size and the position coordinates of the origin and destination.
CoT: This prompt guides the model through a structured planning procedure. It instructs the LLM to first validate that the start and end points are not blocked, identify valid moves, and iteratively construct the path. This design encourages the model to explicitly reason about each step of the planning process.
FewShot-Base: It presents two examples with complete input-output pairs, including obstacle descriptions and optimal paths, allowing the model to learn the structure of the task by analogy.
4.2.2 Algorithms of Thought (AoT) Prompt
To evaluate the capability of LLMs in applying classical algorithms, we designed the Algorithms of Thought (AoT) prompting framework. This framework guides LLMs to mimic classical path planning algorithms by prompting them to apply algorithmic reasoning, analyze the task environment, constrain movement directions, explain algorithmic principles, and adhere to a clear output format.
AoT-DFS: This prompt is designed to guide the model in following the Depth-First Search (DFS).Instead of outlining the algorithm formally, the prompt incorporates its core mechanisms—ordered expansion, backtracking, and visit marking—into the structure of the instruction itself. By structuring the order of exploration and discouraging revisits,it leads the model to follow a behavior pattern similar to DFS.
AoT-A*: This prompt adapts the logic of the A* algorithm to guide the model in path planning. It introduces key elements of A*, such as cost estimation, heuristic evaluation, and goal-directed prioritization. The prompt defines valid moves, explains how to measure distance using the Manhattan heuristic, and describes how to compare path options. The model is prompted to evaluate each possible move based on its distance to the goal and the cost incurred so far, simulating how A* selects efficient paths.
AoT-Dijkstra: It provides a structured framework by embedding Dijkstra’s core logic – greedy distance minimization, neighbor evaluation, and path optimality. In detail, it enforces a multi-step workflow: initializing Dijkstra’s algorithm (infinite distances except the starting cell), iteratively updating route costs via priority queue simulation, and reconstructing the shortest route through parent-node backtracking.
4.2.3 AoT with Example (Algo-Shot) Prompt
To evaluate the impact of integrating LLMs with traditional algorithms under different shot strategies, we design two distinct shot prompt configurations as follows:
Algo-Direct: It integrates the AoT-Dijkstra prompt with the same example used in FewShot-Base - providing two cases in the grid, with each case including only the start point, end point, obstacles, and the optimal path computed using the Dijkstra algorithm.
Algo-Reasoning: Different from Algo-Direct, Algo-Reasoning combines AoT-Dijkstra with example that provides a detailed demonstration of the path planning reasoning steps from the starting point to the destination. To align with the reasoning process in this example, we configured the internal steps of AoT-Dijkstra in greater detail, clearly specifying how various situations encountered during path planning are handled.
5 Experimental Analysis
In this section, building upon the four core research questions (RQ) introduced in the Introduction, we conduct the following analyses to provide deeper insights into how various prompting strategies interact with traditional algorithmic guidance under varying conditions.
RQ1: How does the Algorithms of Thought (AoT) prompt perform compared to the Independent Route Planning prompt?
In Figure 3, we find that AoT prompting methods significantly outperform the Vanilla prompt across all metrics except AoT-DFS, especially in more complex environments. Moreover, the integration of LLMs with optimal pathfinding algorithms (e.g., Dijkstra or A*) achieves better planning performance compared to their integration with uninformed search methods like DFS. These methods show clear advantages in terms of FR and OR, indicating stronger path planning capabilities. However, a notable trade-off exists for AoT-DFS compared to Vanilla: despite the improvements in FR and OR, GM and MSE can sometimes be higher.
Compared to the more advanced independent prompting strategy CoT, AoT achieves comparable or even superior performance, particularly on Qwen2.5-7B, as illustrated in Figure 13. Figure 5 further highlights that AoT surpasses CoT in terms of geometric mean (GM), reflecting higher overall route quality. However, CoT exhibits more consistent performance across all models, with fewer failure cases as shown in Figure 2. In contrast, within the AoT framework, only AoT-Dijkstra consistently maintains robust results. These findings suggest that CoT’s step-by-step reasoning offers greater reliability across model sizes, while AoT’s effectiveness depends more on the model’s ability to accurately internalize and apply classical algorithmic logic.
FewShot-Base, despite its simplicity compared to advanced prompting strategies such as CoT and AoT, achieves performance comparable to AoT-Dijkstra by relying on minimal task demonstrations involving only obstacles, start/end points, and optimal paths. This highlights the efficacy of lightweight imitation-based learning and illustrates the strong transfer and generalization capabilities of LLMs in path planning tasks.
Overall, CoT demonstrates the lowest total error count, while AoT prompts shows comparable performance overall. However, AoT exhibits significant start/end mismatch issues in smaller models (e.g., Qwen2.5-7B).
RQ2: How few shots influence LLM’s planning ability when guided by traditional algorithms?
Our experiments comparing AoT-Dijkstra with different shot strategies (Figure 6) reveal that the Algo-Direct prompt generally enhances AoT-Dijkstra performance in most metrics, while the Algo-Reasoning setting improves its performance specifically in GM and MSE at the cost of reduced FR and OR. However, due to the limited capabilities of Qwen2.5-7B, Algo-Reasoning could not produce valid outputs on this smaller model. When using the same example, Algo-Direct prompting outperforms FewShot-Base by using the AoT-Dijkstra prompt while preserving example demonstration patterns.
RQ3: How does the size of the model influence the traditional algorithm-guided prompt?
Figure 2 demonstrates that within the Qwen2.5 model series, scaling from 7B to 32B leads to substantial improvements in FR and RT. The 72B model further enhances FR while significantly boosting OR and reducing GM and MSE. This suggests that the FR gains observed in Qwen2.5-32B come at the cost of higher GM and MSE. Notably, Qwen2.5-72B achieves performance comparable to models such as GPT-4 Turbo(1 trillion) and DeepSeek-V3(671B), suggesting that simply increasing model size in GridRoute tasks eventually encounters a scaling law bottleneck. Additionally, as model size increases, errors such as illegal moves and start-end mismatches become significantly less frequent and are increasingly concentrated in path-through-obstacle cases. In most prompting strategies, inference speed tends to improve with larger model sizes, especially under AoT-Dijkstra and vanilla prompts. However, AoT-DFS shows less consistent improvement and even experiences relatively increased latency from 7B to 32B, indicating that response time is not strictly correlated with model size.
RQ4: How does map size influence traditional algorithm-guided prompt?
As shown in Figure 7, the scaling of map size is a critical factor affecting the path planning performance across all evaluation metrics. As the map size increases, both the FR and OR exhibit a systematic decline, while the RT correspondingly increases. This scaling effect holds consistently regardless of model size, although the influence of model size diminishes as the map becomes larger—for instance, on a 30×30 grid, the performance of Qwen2.5-32B approaches that of Qwen2.5-72B. Notably, the nature of errors evolves with increasing map complexity, transitioning from primarily " Path though obstacles" to more intricate types of errors. While model size and architectural differences do influence performance, their impact is secondary to the fundamental constraints imposed by map dimensionality. In particular, the CoT prompting strategy exhibits more generalized adaptability across varying map sizes compared to the AoT approach. This suggests that with appropriate prompting strategies, LLMs can achieve path planning performance comparable to or even competitive with AoT methods.
6 Conclusion
We introduced GridRoute, a benchmark for systematically evaluating LLM-based route planning with guidance from classical algorithms. Our proposed AoT prompting framework enables LLMs to incorporate structured reasoning inspired by A*, Dijkstra, and DFS, significantly improving performance compared with Vallina, especially in complex environments. AoT performs comparably to or better than CoT, while example-based AoT variants further enhance generalization. Although CoT offers more stable performance across model sizes, our findings highlight the complementary strengths of symbolic algorithms and LLMs, suggesting promising directions for hybrid neuro-symbolic planning systems.
7 Limitations
Our current benchmark shows that the performance of 32B and 70B models tends to converge as map dimensions increase. However, the performance limits of LLMs on ultra-large-scale maps remain unclear and require extended experiments to assess the trade-off between computational efficiency and planning accuracy. Notably, while increasing model size and refining prompting strategies significantly reduce most types of error, critical issues such as path traversal through obstacles persist. These could potentially be mitigated through mechanisms like obstacle collision backtracking or targeted fine-tuning. Furthermore, existing studies focus solely on single-goal planning in open maps. The GridRoute framework, however, can be naturally extended to multi-goal cooperative planning scenarios—such as logistics delivery or multi-task inspection—marking a promising direction for hybrid decision-making that integrates traditional algorithms with LLMs.
Appendix A Prompt Design
A.1 Independent Route Planning Prompts
Given these obstacle coordinates in a grid{obstacles}: Please independently plan a continuous path of coordinates from the starting point ({start_x}, {start_y})to the ending point ({end_x}, {end_y}), avoiding all obstacles listed above. The path must follow these rules: 1. Every point must differ by exactly 1 in either x or y direction from the previous point (no diagonal moves allowed). 2. The output should be in the format: [(x1, y1), (x2, y2), (x3, y3), …]. 3. If a valid path cannot be created without touching an obstacle, return an empty path []. Return only the path without any additional explanation.
Here are the obstacle coordinates in the grid:{obstacles}" You need to plan a path from the starting point ({start_x}, {start_y}) to the ending point ({end_x, end_y) while avoiding the obstacles. To ensure accuracy, follow these steps step by step: 1. Verify that the starting point {(start_x},{start_y)} and ending point ({end_x, {end_y}) are not obstacles. 2. List all valid moves for a point in the grid. A move is valid if: - It differs by exactly 1 in either the x or y direction from the previous point. - It does not overlap with any obstacle. 3. Starting from ({start_x}, {start_y}), iteratively choose the next valid move to build a path to ({end_x}, {end_y}). 4. If there is no valid path, return an empty path []. Return the final path in this format: [(x1, y1), (x2, y2), (x3, y3), …]. If there are any errors or no valid paths, return []. Make sure to provide only the path in your response without any extra explanations.
Plan the shortest path from the starting point ({start_x}, {start_y}) to the endpoint ({end_x}, {end_y}) while avoiding rectangular obstacles defined as {obstacles}. Follow these steps: 1. Analyze the grid to identify the start, end, and rectangular obstacle locations. 2. Each rectangle is defined by its top-left corner (x1, y1) and bottom-right corner (x2, y2). A grid cell (x, y) is an obstacle if and . 3. Define valid moves: -Right: -Left: -Down: -Up: 4. Build the path: - Start at ({start_x}, {start_y}). - At each step, evaluate valid moves and select the one that minimizes the distance to ({end_x}, {end_y}). 5. If no valid moves are available, return an empty path []. 6. Output the result in this format: [(x1, y1), (x2, y2), …, (xn, yn)]. Example 1: Start: (3, 7), End: (4, 3), Obstacles: [((1, 2), (3, 4)), ((2, 5), (4, 6))] Path: [(3, 7), (4, 7), (5, 7), (5, 6), (5, 5), (5, 4), (5, 3), (4, 3)] Example 2: Start: (2, 4), End: (7, 5), Obstacles: [((3, 4), (5, 6))] Path: [(2, 4), (2, 3), (3, 3), (4, 3), (5, 3), (6, 3), (7, 3), (7, 4), (7, 5)] Return only the path as output, without additional explanations.
A.2 Algorithm of Thought(AoT) Prompts
Plan a path from the starting point ({start_x}, {start_y}) to the endpoint ({end_x}, {end_y}) while avoiding rectangular obstacles defined as {obstacles}. Use DFS algorithm to explore possible paths. Follow these steps: 1. Analyze the grid to identify the start, end, and rectangular obstacle locations. 2. Each rectangle is defined by its top-left corner (x1, y1) and bottom-right corner (x2, y2). A grid cell (x, y) is an obstacle if and . 3. Define valid moves (same priority order): - Right: (x, y+1) - Left: (x, y-1) - Down: (x+1, y) - Up: (x-1, y) - All moves must stay within the grid bounds and avoid obstacle cells. 4. Use Depth-First Search (DFS) algorithm: - Initialize a stack with the start cell and its path [(start_x, start_y)] - Create a visited set to track explored cells - While stack is not empty: - Pop the last cell and its path from the stack - If current cell is the endpoint, return the complete path - If cell not in visited: - Mark cell as visited - Push all valid neighboring cells to the stack (FILO order): * For each neighbor in move order (Right→Left→Down→Up): - If neighbor is within grid bounds - If neighbor is not an obstacle - If neighbor not in visited - Push (neighbor_coords, path + [neighbor_coords]) to stack 5. Termination conditions: - If stack becomes empty before reaching endpoint, return empty path []. - First valid path found using DFS should be returned. 6. Output the result in this format: [(x1, y1), (x2, y2), …, (xn, yn)]. Return only the path as output, without additional explanations.
Plan the shortest path from the starting point ({start_x}, {start_y}) to the endpoint ({end_x}, {end_y}) while avoiding rectangular obstacles defined as {obstacles}. Use A* algorithm to calculate the shortest path. Follow these steps: 1. Analyze the grid to identify the start, end, and rectangular obstacle locations. 2. Each rectangle is defined by its top-left corner (x1, y1) and bottom-right corner (x2, y2). A grid cell (x, y) is an obstacle if and . 3. Define valid moves: - Right: (x, y+1) - Left: (x, y-1) - Down: (x+1, y) - Up: (x-1, y) - All moves must stay within the grid bounds and avoid obstacle cells. 4. Use A* algorithm: - Initialize g_score (cost from start) for all cells to infinity, except the start cell which is 0. - Initialize f_score (g_score + heuristic) for all cells to infinity, except the start cell which is heuristic(start, end). - Use a priority queue to repeatedly select the grid cell with the lowest f_score. - For each selected cell, evaluate its valid neighbors: - Calculate the tentative g_score for the neighbor through the current cell. - If the tentative g_score is lower than the neighbor’s current g_score, update the neighbor’s parent, g_score, and f_score (g_score + heuristic(neighbor, end)). - Add the neighbor to the priority queue if it’s not already there or if this path is better. - Use the Manhattan distance as the heuristic function: h((x1, y1), (x2, y2)) = abs(x1 - x2) + abs(y1 - y2). - Continue until the endpoint (end_x, end_y) is reached or the priority queue is empty. 5. Construct the path: - Start at (start_x, start_y). - Trace back from the endpoint to the starting point using the recorded parent cells to reconstruct the shortest path. - If no valid moves are available, return an empty path []. 6. Output the result in this format: [(x1, y1), (x2, y2), …, (xn, yn)]. Return only the path as output, without additional explanations.
"Plan the shortest path from the starting point ({start_x}, {start_y}) to the endpoint ({end_x}, {end_y}) while avoiding rectangular obstacles defined as {obstacles}. Use Dijkstra’s algorithm to calculate the shortest path. Follow these steps:" 1. Analyze the grid to identify the start, end, and rectangular obstacle locations. 2. Each rectangle is defined by its top-left corner (x1, y1) and bottom-right corner (x2, y2). A grid cell (x, y) is an obstacle if and . 3. Define valid moves: - Right: (x, y+1) - Left: (x, y-1) - Down: (x+1, y) - Up: (x-1, y) - All moves must stay within the grid bounds and avoid obstacle cells. 4. Use Dijkstra’s algorithm: - Assign an initial distance of infinity to all grid cells except the starting point, which should have a distance of 0. - Use a priority queue to repeatedly select the grid cell with the smallest tentative distance. - For each selected cell, evaluate its valid neighbors and update their tentative distances if a shorter path is found. - Continue until the endpoint ({end_x}, {end_y}) is reached or all possible paths have been explored. 5. Construct the path:" - Start at ({start_x}, {start_y}). - Trace back from the endpoint to the starting point using the recorded parent cells to reconstruct the shortest path. - If no valid moves are available, return an empty path []. 6. Output the result in this format: [(x1, y1), (x2, y2), …, (xn, yn)]. Return only the path as output, without additional explanations.
A.3 AoT with Example (Algo-Shot) Prompt
"Plan the shortest path from the starting point ({start_x}, {start_y}) to the endpoint ({end_x}, {end_y}) while avoiding rectangular obstacles defined as {obstacles}. Use Dijkstra’s algorithm to calculate the shortest path. Follow these steps:" 1. Analyze the grid to identify the start, end, and rectangular obstacle locations. 2. Each rectangle is defined by its top-left corner (x1, y1) and bottom-right corner (x2, y2). A grid cell (x, y) is an obstacle if and . 3. Define valid moves: - Right: (x, y+1) - Left: (x, y-1) - Down: (x+1, y) - Up: (x-1, y) - All moves must stay within the grid bounds and avoid obstacle cells. 4. Use Dijkstra’s algorithm: - Assign an initial distance of infinity to all grid cells except the starting point, which should have a distance of 0. - Use a priority queue to repeatedly select the grid cell with the smallest tentative distance. - For each selected cell, evaluate its valid neighbors and update their tentative distances if a shorter path is found. - Continue until the endpoint ({end_x}, {end_y}) is reached or all possible paths have been explored. 5. Construct the path:" - Start at ({start_x}, {start_y}). - Trace back from the endpoint to the starting point using the recorded parent cells to reconstruct the shortest path. - If no valid moves are available, return an empty path []. 6. Output the result in this format: [(x1, y1), (x2, y2), …, (xn, yn)]. Example 1: Start: (3, 7), End: (4, 3), Obstacles: [((1, 2), (3, 4)), ((2, 5), (4, 6))] Path: [(3, 7), (4, 7), (5, 7), (5, 6), (5, 5), (5, 4), (5, 3), (4, 3)] Example 2: Start: (2, 4), End: (7, 5), Obstacles: [((3, 4), (5, 6))] Path: [(2, 4), (2, 3), (3, 3), (4, 3), (5, 3), (6, 3), (7, 3), (7, 4), (7, 5)] Return only the path as output, without additional explanations.
Here are the obstacle coordinates in the grid:{obstacles} You need to plan a path from the starting point ({start_x}, {start_y}) to the ending point ({end_x}, {end_y}) while avoiding the obstacles. Use Dijkstra’s algorithm principles to ensure the path is valid and follows the shortest possible route. Follow these steps: 1. Verify the starting point ({start_x}, {start_y}) and ending point ({end_x}, {end_y}) are not obstacles. 2. Define the cost of reaching a point in the grid: - Start by assigning a cost of 0 to the starting point ({start_x}, {start_y}). - For all other points, initialize the cost as infinity (). 3. Use a priority queue to iteratively explore points in the grid: - Begin with the starting point ({start_x}, {start_y}) in the queue. - At each step, select the point with the lowest current cost. - For each valid move (up, down, left, right): a. If moving to a neighboring point reduces its cost, update the cost and add the point to the queue. b. A move is valid if it does not overlap with any obstacle and stays within the grid. 4. Repeat this process until reaching the ending point ({end_x}, {end_y}) or exhausting all valid points in the queue: - If the ending point is reached, reconstruct the path by tracing back from the endpoint to the start using the cost values. - If no valid path exists, return an empty path []. 5. Return the final path in this format: [(x1, y1), (x2, y2), (x3, y3), …]. If there are any errors or no valid paths, return []. Example: Start: (3, 7), End: (4, 3), Obstacles: [((1, 2), (3, 4)), ((2, 5), (4, 6))], grid: 10 Evaluate valid neighbors in given map: - Starting at (3, 7), distance is 0 : - (3, 6): Obstructed, skip. - (4, 7): Valid, update distance to 1. - (2, 7): Valid, update distance to 1. - (3, 8): Valid, update distance to 1. - To (3, 6): Obstructed point. - To (3, 8), recorded distance is 1: - (3, 7): Start point, distance is 0. - (4, 8): Valid, update distance to 2. - (2, 8): Valid, update distance to 2. - (3, 9): Valid, update distance to 2. - To (3, 9), recorded distance is 2: - (3, 8): Already visited, distance is,3, but recorded distance is , so distance stays at 1. - (4, 9): Valid, update distance to 3. - (2, 9): Valid, update distance to 3. - (3, 10): Out of grids, skip. - To (4, 9), recorded distance is 3: - (4, 8): Already visited, distance is 4, but recorded distance is , so distance stays at 2. - (5, 9): Valid, update distance to 4. - (3, 9): Already visited, distance is 4, but recorded distance is , so distance stays at 2. - (4, 10): Out of grids, skip - To (5, 9), recorded distance is 4: - (5, 8): Valid, update distance to 5. - (6, 9): Valid, update distance to 5. - (4, 9): Already visited, distance is 5, but recorded distance is , so distance stays at 3. - (5, 10): Out of grids, skip. - To (4, 8), recorded distance is 2: - (4, 7): Already visited, distance is 3, but recorded distance is 1 < 3, so distance stays at 1. - (5, 8): Already visited, distance is 3, but recorded distance is 5 > 3, so distance updates to 3. - (3, 8): Already visited, distance is 3, but recorded distance is 1 < 3, so distance stays at 1. - (4, 9): Already visited, distance is 3 which is equal to recorded distance, so distance stays at 3. - Traverse all point in the map - Result: the distance of (4,3) is 7, which means (3,7) to (4,3) need 7 steps. - Generated Path: [(3, 7), (4, 7), (5, 7), (5, 6), (5, 5), (5, 4), (4, 4), (4, 3)] Now, apply the same logic to plan a path from ({start_x}, {start_y}) to ({end_x}, {end_y}) while avoiding the obstacles: Make sure to provide only the path in your response without any extra explanations.
Appendix B Results
| Model | Prompt | 10 size | 20 size | 30 size | |||||||||||||||
| CR | FR | OR | GM | MSE | RT(s) | CR | FR | OR | GM | MSE | RT(s) | CR | FR | OR | GM | MSE | RT(s) | ||
| deepseek-V3 | Vanilla | 67 | 48 | 47 | 104.1 | 3 | 1.9 | 37 | 26 | 26 | 100 | 0 | 2.4 | 19 | 11 | 10 | 109.2 | 29.5 | 3.0 |
| qwen2.5-7b-instruct | 100 | 48 | 47 | 101.5 | 8.3 | 2.4 | 99 | 25 | 24 | 101.4 | 1.4 | 12.5 | 99 | 13 | 13 | 100 | 0 | 35.1 | |
| qwen2.5-32b-instruct | 98 | 63 | 45 | 121.1 | 12.9 | 2.2 | 99 | 52 | 48 | 104.8 | 4.9 | 4.3 | 98 | 64 | 61 | 104.4 | 17.4 | 7.0 | |
| qwen2.5-72b-instruct | 100 | 72 | 69 | 104.0 | 3.3 | 1.9 | 100 | 61 | 58 | 105.3 | 16.7 | 3.7 | 100 | 57 | 54 | 105.4 | 29.8 | 5.9 | |
| llama3-70b-instruct | 100 | 69 | 52 | 124.5 | 15.2 | 1.2 | 100 | 57 | 48 | 112.5 | 35.1 | 1.9 | 100 | 54 | 48 | 113.9 | 93.3 | 2.8 | |
| ChatGPT-4 | 100 | 76 | 63 | 112.2 | 11.5 | 2.3 | 99 | 61 | 48 | 113.5 | 36.8 | 3.7 | 98 | 62 | 52 | 112.2 | 101.0 | 5.1 | |
| deepseek-V3 | CoT | 100 | 73 | 72 | 100.39 | 0.054 | 2.95 | 93 | 68 | 68 | 100.00 | 0 | 5.54 | 76 | 53 | 53 | 100.00 | 0 | 7.54 |
| qwen2.5-7b-instruct | 100 | 64 | 58 | 106.12 | 2.18 | 1.89 | 100 | 66 | 64 | 101.57 | 1.03 | 3.94 | 93 | 54 | 50 | 104.26 | 12.96 | 6.01 | |
| qwen2.5-32b-instruct | 100 | 63 | 57 | 105.54 | 2.03 | 2.21 | 100 | 67 | 66 | 100.50 | 0.05 | 4.03 | 94 | 54 | 51 | 103.03 | 6.96 | 6.27 | |
| qwen2.5-72b-instruct | 100 | 64 | 58 | 106.12 | 2.18 | 1.89 | 100 | 66 | 64 | 101.57 | 1.03 | 3.94 | 93 | 54 | 50 | 104.26 | 12.96 | 6.01 | |
| llama3-70b-instruct | 100 | 70 | 64 | 105.86 | 2.91 | 1.24 | 100 | 60 | 60 | 100 | 0 | 1.87 | 100 | 62 | 62 | 100 | 0 | 2.32 | |
| ChatGPT-4 | 100 | 84 | 77 | 102.31 | 0.71 | 2.47 | 100 | 75 | 64 | 104.79 | 21.33 | 4.00 | 100 | 69 | 64 | 103 | 39.13 | 6.25 | |
| deepseek-V3 | Few-shot | 100 | 85 | 84 | 100.03 | 0.18 | 1.94 | 100 | 70 | 70 | 100 | 0 | 2.61 | 100 | 67 | 67 | 100 | 0 | 3.37 |
| qwen2.5-7b-instruct | Learning | 100 | 44 | 38 | 109.28 | 10.09 | 3.47 | 100 | 37 | 35 | 102.54 | 4.32 | 4.48 | 99 | 21 | 19 | 102.89 | 4.76 | 8.71 |
| qwen2.5-32b-instruct | 100 | 65 | 55 | 108.18 | 3.81 | 2.05 | 100 | 54 | 46 | 108.67 | 26.37 | 4.62 | 100 | 60 | 42 | 121.79 | 134.03 | 6.85 | |
| qwen2.5-72b-instruct | 100 | 76 | 71 | 102.85 | 2.10 | 2.15 | 100 | 62 | 60 | 101.57 | 2.38 | 4.30 | 100 | 60 | 59 | 100.24 | 106.67 | 4.30 | |
| llama3-70b-instruct | 100 | 70 | 68 | 102.32 | 0.28 | 1.06 | 99 | 56 | 56 | 100 | 0 | 11.57 | 100 | 56 | 56 | 100 | 0 | 9.90 | |
| ChatGPT-4 | 85 | 62 | 62 | 100 | 0 | 2.9 | 98 | 68 | 65 | 101.1 | 1.23 | 4.26 | 93 | 67 | 62 | 100.3 | 38.02 | 6.88 | |
| Model | Prompt | 10 size | 20 size | 30 size | |||||||||||||||
| CR | FR | OR | GM | MSE | RT(s) | CR | FR | OR | GM | MSE | RT(s) | CR | FR | OR | GM | MSE | RT(s) | ||
| deepseek-V3 | AoT | 100 | 79 | 65 | 111.64 | 5.87 | 8.38 | 100 | 67 | 62 | 105.28 | 25.43 | 10.26 | 100 | 72 | 59 | 114.34 | 131.77 | 14.40 |
| qwen2.5-7b-instruct | (DFS) | 100 | 51 | 47 | 105.68 | 2.11 | 2.59 | 100 | 27 | 26 | 101.19 | 2.37 | 3.88 | 100 | 19 | 15 | 106.26 | 9.89 | 12.39 |
| qwen2.5-32b-instruct | 100 | 62 | 49 | 116.14 | 4.77 | 3.09 | 100 | 64 | 50 | 118.79 | 23.375 | 4.90 | 97 | 52 | 43 | 112.98 | 112.85 | 8.00 | |
| qwen2.5-72b-instruct | 100 | 77 | 69 | 104.40 | 1.82 | 3.14 | 100 | 62 | 58 | 101.83 | 0.83 | 11.69 | 100 | 60 | 51 | 107.20 | 25.13 | 11.52 | |
| llama3.1-70b-instruct | 93 | 61 | 52 | 115.13 | 6.22 | 12.55 | 100 | 54 | 54 | 100 | 0 | 3.64 | 91 | 57 | 53 | 109.21 | 80.49 | 10.50 | |
| ChatGPT-4 | 3 | 3 | 3 | 100 | 0 | 5.04 | 1 | 1 | 1 | 100 | 0 | 11.85 | 1 | 1 | 1 | 100 | 0 | 5.62 | |
| deepseek-V3 | AoT | 100 | 80 | 80 | 100 | 0 | 3.27 | 100 | 74 | 74 | 100 | 0 | 7.13 | 100 | 66 | 66 | 100 | 0 | 6.43 |
| qwen2.5-7b-instruct | (A*) | 100 | 57 | 55 | 101.16 | 1.19 | 3.03 | 100 | 36 | 34 | 102.97 | 8 | 7.85 | 99 | 25 | 24 | 104.49 | 36 | 7.13 |
| qwen2.5-32b-instruct | 99 | 67 | 59 | 106.61 | 1.31 | 2.21 | 100 | 61 | 59 | 100.94 | 2.22 | 3.80 | 100 | 54 | 50 | 105.38 | 37.18 | 5.49 | |
| qwen2.5-72b-instruct | 99 | 77 | 73 | 101.57 | 0.36 | 1.89 | 100 | 62 | 62 | 100 | 0 | 3.00 | 100 | 61 | 60 | 100.23 | 1.04 | 4.89 | |
| llama3.1-70b-instruct | 100 | 82 | 81 | 100.30 | 0.05 | 2.62 | 100 | 74 | 73 | 100.63 | 1.94 | 4.18 | 100 | 70 | 67 | 102.06 | 29.77 | 10.02 | |
| ChatGPT-4 | 84 | 65 | 61 | 102.26 | 1.16 | 4.90 | 66 | 48 | 45 | 100.68 | 0.25 | 10.35 | 54 | 41 | 39 | 101.39 | 19.21 | 10.10 | |
| deepseek-V3 | AoT | 100 | 81 | 79 | 101.40 | 0.49 | 1.89 | 100 | 69 | 69 | 100 | 0 | 2.57 | 100 | 73 | 72 | 101.13 | 17.75 | 3.46 |
| qwen2.5-7b-instruct | (Dijkstra) | 100 | 57 | 55 | 1.0094 | 0.35 | 2.00 | 100 | 33 | 31 | 101.88 | 4.48 | 2.87 | 100 | 23 | 20 | 106.38 | 17.04 | 66.42 |
| qwen2.5-32b-instruct | 100 | 64 | 55 | 107.63 | 3.625 | 1.79 | 100 | 0.51 | 0.45 | 106.52 | 12.08 | 3.77 | 99 | 59 | 42 | 120.89 | 114.92 | 5.79 | |
| qwen2.5-72b-instruct | 100 | 78 | 75 | 101.05 | 0.56 | 2.09 | 100 | 62 | 61 | 100.54 | 0.26 | 3.98 | 1 | 66 | 66 | 1 | 0 | 5.89 | |
| llama3.1-70b-instruct | 96 | 67 | 64 | 103.44 | 1.25 | 1.50 | 95 | 52 | 51 | 101.55 | 4.92 | 2.85 | 98 | 63 | 61 | 103.81 | 14.29 | 4.34 | |
| ChatGPT-4 | 99 | 78 | 76 | 100.72 | 0.10 | 2.99 | 95 | 66 | 59 | 101.88 | 0.79 | 5.82 | 76 | 50 | 47 | 103.21 | 42.88 | 7.65 | |
| Model | Prompt | 10 size | 20 size | 30 size | |||||||||||||||
| CR | FR | OR | GM | MSE | RT(s) | CR | FR | OR | GM | MSE | RT(s) | CR | FR | OR | GM | MSE | RT(s) | ||
| deepseek-V3 | Easy | 100 | 84 | 82 | 100.54 | 0.10 | 4.79 | 100 | 76 | 74 | 101.61 | 5.26 | 6.22 | 100 | 71 | 70 | 101.56 | 3.61 | 8.29 |
| qwen2.5-7b-instruct | shot | 100 | 50 | 43 | 109.82 | 3.12 | 2.10 | 100 | 40 | 36 | 103.06 | 3.1 | 3.47 | 97 | 31 | 25 | 105.40 | 14.97 | 7.24 |
| qwen2.5-32b-instruct | 100 | 71 | 58 | 109.22 | 3.83 | 2.04 | 99 | 54 | 49 | 106.75 | 22.44 | 3.67 | 98 | 59 | 51 | 105.12 | 19.53 | 5.00 | |
| qwen2.5-72b-instruct | 100 | 77 | 71 | 102.48 | 0.31 | 2.18 | 86 | 52 | 50 | 102.89 | 2.85 | 4.09 | 80 | 53 | 53 | 100 | 0 | 4.97 | |
| llama3.1-70b-instruct | 98 | 74 | 74 | 100 | 0 | 4.65 | 99 | 64 | 63 | 101.24 | 0.56 | 4.26 | 97 | 60 | 59 | 102.34 | 9.6 | 15.51 | |
| ChatGPT-4 | 85 | 69 | 69 | 100 | 0 | 10.89 | 74 | 49 | 46 | 102.22 | 10.29 | 13.74 | 62 | 44 | 44 | 100 | 0 | 17.47 | |
| deepseek-V3 | Complex | 100 | 82 | 79 | 100.82 | 0.15 | 2.06 | 100 | 67 | 67 | 100 | 0 | 2.91 | 100 | 64 | 62 | 100.96 | 3.81 | 4.06 |
| qwen2.5-7b-instruct | shot | 0 | 0 | 0 | NaN | NaN | NaN | 0 | 0 | 0 | NaN | NaN | NaN | 0 | 0 | 0 | NaN | NaN | NaN |
| qwen2.5-32b-instruct | 95 | 39 | 33 | 110.15 | 2.77 | 2.65 | 93 | 49 | 41 | 110.00 | 19.27 | 5.78 | 98 | 39 | 0.3 | 112.20 | 72.51 | 7.09 | |
| qwen2.5-72b-instruct | 92 | 66 | 56 | 108.21 | 2.24 | 2.16 | 97 | 62 | 62 | 100 | 0 | 3.68 | 99 | 56 | 54 | 100.46 | 1.42 | 5.64 | |
| llama3.1-70b-instruct | 100 | 72 | 69 | 102.59 | 0.33 | 1.57 | 100 | 62 | 59 | 103.62 | 15.61 | 1.96 | 100 | 56 | 55 | 100.64 | 2.57 | 11.69 | |
| ChatGPT-4 | 100 | 74 | 71 | 101.62 | 0.76 | 2.31 | 100 | 74 | 71 | 100.79 | 0.16 | 3.07 | 99 | 67 | 65 | 101.61 | 20.84 | 4.61 | |